OpenStudy (ksaimouli):

a circle is tangent to the y-axis at y=3 and has one x-intercept ast x=1(a) determine the other x-intercept

OpenStudy (anonymous):



OpenStudy (anonymous):

hmm.... one way i can see you do this is solve for the center, (x, y) by using the distance formula: \(\large \sqrt{(x-0)^2+(y-3)^2}=\sqrt{(x-1)^2+(y-0)^2} \) solve this equation for x and y ... then, since the radius is the x coordinate of the center, you can write down the equation for the circle and therefore find the other x-intercept.

OpenStudy (anonymous):

yep... this is a multi-step problem....

OpenStudy (ksaimouli):

so i can use any one distance formula which u have given right